OpenStudy (anonymous):

Let f(x) =√(x+1.) compute the Taylor series around a=3 for the first three terms. Use this to estimate √(4.1.) What is the error?

OpenStudy (anonymous):

\[2+\frac{x-3}{4}-\frac{1}{64} (x-3)^2 \] The series evaluated at 3.1 is 2.02484 \[\sqrt{4.1} = 2.02485\] \[\text{Error}: 1*10^{-5} \] Here is the series with an extra term if you need it. \[2+\frac{x-3}{4}-\frac{1}{64} (x-3)^2+\frac{1}{512} (x-3)^3 \]